OpenStudy (anonymous):

Use the derivative of the function y - f(x) to find the points at which f has a a) local max b) local min c) point of inflection

OpenStudy (anonymous):

y'= (x-1)^2 (x-2)

OpenStudy (anonymous):

equate the derivative to 0 to find any max or min equate the second derivative to 0 to find any poi

OpenStudy (anonymous):

ohhh ok thanks!@

OpenStudy (anonymous):

so it's (x-1)^2 (x-2) = 0 x = 2, 0 and does (x-1)^2 has an x value?

OpenStudy (anonymous):

of course it does

OpenStudy (anonymous):

oh so it's -1 and 1 also?

OpenStudy (anonymous):

(x-1)^2 = 0 solve for x

OpenStudy (anonymous):

(x-1)(x-1)(x-2) = 0 correct? to find max and mins?

OpenStudy (anonymous):

since (x-1)^2 is the same as (x-1) (x-1)

OpenStudy (anonymous):

^ no (x^2 -1) = (x-1)(x+1)

OpenStudy (anonymous):

how is it x^2 - 1? if the squared is outside the parenthesis?? it's not (x^2-1) because it says (x-1)^2? isn't that different? :0

OpenStudy (anonymous):

what i'm saying is (x-1)^2 =/= (x-1)(x+1)

OpenStudy (anonymous):

does not equal?